# Dirichlet-to-Neumann operator for the Helmholtz problem with general wavenumbers on the $n$-sphere

### v1

@@ Abstract
This paper considers the Helmholtz problem in the exterior of a ball with Dirichlet boundary conditions and radiation conditions imposed at infinity. The differential Helmholtz operator depends on the complex wavenumber with non-negative real part and is formulated for general spatial dimension. We prove wavenumber explicit continuity estimates of the corresponding Dirichlet-to-Neumann ($\operatorname*{DtN}$) operator which are valid for all wavenumbers under consideration and do not deteriorate as they tend to zero.
@@ 1 Introduction
The theoretical justification and the practical implementation of this approach rely on a thorough understanding of the involved DtN operator. In particular, the standard proofs of the well-posedenss of the truncated problem require definiteness properties and the continuity of the DtN operator [7, 8, 13, 17]. However, such properties have only been established for purely imaginary wavenumbers in two and three [2, 13, 14] dimensions based on explicit Fourier expansion, preventing a unified theory for general wavenumbers. To close this gap, this paper presents a unified mathematical analysis of spherical DtN operators for Helmholtz problems with general wavenumbers in any spatial dimension and provides a clearer mathematical understanding of the important definiteness characteristics of the DtN operator. The new insights also enable a fine-tuning of control parameters in numerical methods depending on the real and imaginary part of the wavenumber such as [1] in future research.
@@
a) The wavenumber-dependent DtN operator admits an explicit series representation in terms of eigenfunctions of the Laplace-Beltrami operator and Bessel functions. A novel wavenumber-explicit control of the series coefficients enables sharp bounds on the real and imaginary part of the DtN operator for general wavenumbers and spatial dimensions. The results extend the existing theory for purely imaginary wavenumbers, refining the bounds in [13, Sec. 3] for two dimensions and recovering the classical results [14, Eqn. (2.6.23)–(2.6.24)] for three dimensions.
b) The solution space of the Helmholtz equation in the truncated domain with transparent ($\operatorname*{DtN}$) boundary conditions on the artificial boundary satisfies a novel Friedrichs-type inequality. This enables wavenumber-explicit Sobolev and trace estimates which do not deteriorate for small wavenumbers and are valid for all complex wavenumbers with non-negative real part.
@@
Section 3 presents the exterior Helmholtz problem with Dirichlet boundary conditions and discusses the popular choices of the radiation conditions towards infinity. The well-posedness of the resulting exterior problems is established by Proposition 3.2. Its proof follows mainly the theory developed in [12] and is postponed to the appendix. Theorem 3.4 states the important definiteness properties of the DtN operator with sharp bounds of on its real and imaginary parts. The Friedrichs-type inequality described in b) is formulated in Theorem 3.5 and results in Corollary 3.6 on the announced wavenumber-explicit trace estimates.
Section 4 introduces series expansions of the $\operatorname*{DtN}$ operators for the different radiation conditions. Sharp wavenumber-explicit estimates of the series coefficients for non-zero wavenumbers are formulated as Theorem 4.1 and proved in Subsection 4.2. This paper concludes with the the special case of wavenumber zero, i.e., the Laplacian, that requires extra care in Subsection 4.3.
@@ 2 Preliminaries
The physical problem is formulated in $\mathbb{R}^{n}$ for general dimension $n\in\mathopen{}\mathclose{{}\left\{2,3,\ldots}\right\}$. We consider Lipschitz domains $\omega\subset\mathbb{R}^{n}$ which are open and connected (not necessarily bounded) sets whose boundary is compact and locally the graph of a Lipschitz function [12, Def. 3.28]. The classical Lebesgue and Sobolev spaces on $\omega$ are denoted by $L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)$ and $H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ with norms $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}$ and $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)}$. For $\kappa\geq 0$, the space $H_{\operatorname*{loc}}^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ is given by all distributions $v$ on the compactly supported smooth functions $C_{\operatorname*{comp}}^{\infty}\mathopen{}\mathclose{{}\left(\omega}\right)$ such that $\varphi v\in H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ for all $\varphi\in C_{\operatorname*{comp}}^{\infty}\mathopen{}\mathclose{{}\left(\omega}\right)$. For the definition of Sobolev spaces $H^{\kappa}\mathopen{}\mathclose{{}\left(\partial\omega}\right)$ on $\mathopen{}\mathclose{{}\left(n-1}\right)$-dimensional boundaries $\partial\omega$ and their norm $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{H^{\kappa}\mathopen{}\mathclose{{}\left(\Gamma}\right)}$ we refer, e.g., to [12, pp. 96–99]. We define wavenumber-indexed norms [^1] For $s=0$, the quantity $\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right),0}=\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}$ is only a seminorm in $H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)$. for $s\in\mathbb{C}_{\geq 0}$ by
@@
It is well-known that the outer unit normal $n(x)\in\mathbb{S}_{1}$ of the Lipschitz domain $\omega\subset\mathbb{R}^{n}$ exists for almost every $x\in\partial\omega$. The normal derivative $\partial/\partial n:C^{1}\mathopen{}\mathclose{{}\left(\overline{\omega}}\right)\rightarrow C^{0}\mathopen{}\mathclose{{}\left(\Gamma}\right)$ on some relatively open connectivity component $\Gamma\subset\partial\omega$ of the boundary can be continuously extended to a continuous mapping $\gamma_{1}:H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)\rightarrow H^{-1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)$ with
@@
which is equipped by the natural norm
$\mathopen{}\mathclose{{}\left\|\bullet}\right\|_{H^{1}(\omega,\Gamma)}:=\sqrt{\mathopen{}\mathclose{{}\left\|\bullet}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}+\mathopen{}\mathclose{{}\left\|\gamma_{1}\bullet}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Gamma}\right)}^{2}}.$
Let $B_{R}\subset\mathbb{R}^{n}$ denote the open ball of radius $R>0$ about the origin. The outer unit normal vector $x/\mathopen{}\mathclose{{}\left\|x}\right\|$ on the boundary $S_{R}=\partial B_{R}$ points into the unbounded complement $B_{R}^{+}:=\mathbb{R}^{n}\setminus\overline{B_{R}}$ of $B_{R}$ and its derivative in this direction is denoted by $\partial_{r}$. The notation ||$\bullet$|| ∙ | is context-sensitive and may refer to the Lebesgue measure $|\omega|$ of a bounded measurable $n$-dimensional set $\omega\subset\mathbb{R}^{n}$, the surface measure $|\Gamma|$ of an $\mathopen{}\mathclose{{}\left(n-1}\right)$-dimensional manifold $\Gamma\subset\mathbb{R}^{n}$, and the cardinality $|J|$ of a countable set $J$.
@@ 3.1 The Dirichlet Helmholtz problem on the exterior of a ball
$\mathopen{}\mathclose{{}\left\{\begin{array}[c]{rl}\mathopen{}\mathclose{{}\left|u\mathopen{}\mathclose{{}\left(r\xi}\right)}\right|&\leq Cr^{\mu}\\ \mathopen{}\mathclose{{}\left\|\frac{\partial u}{\partial r}\mathopen{}\mathclose{{}\left(r\xi}\right)+su\mathopen{}\mathclose{{}\left(r\xi}\right)}\right\|&\leq Cr^{\mu-1}\end{array}}\right\}\quad\text{as }r\rightarrow\infty\text{ uniformly in }\xi\in S_{1}.$ (3.2a)
@@
In the following proposition we state the well-posedness of these exterior Dirichlet problems. We say that problem (3.1) is well posed for $\operatorname*{IC}_{s}$, if the corresponding solution operator $\operatorname{S_{D}}:H^{1/2}(S_{R})\to H^{1}_{\mathrm{loc}}(B_{R}^{+})$ that assigns each Dirichlet data $g\in H^{1/2}(S_{R})$ the unique solution $\operatorname{S_{D}}g\in H^{1}_{\mathrm{loc}}(B_{R}^{+})$ with $\operatorname{IC}_{s}$ to (3.1), is well-defined and continuous.
@@
      Problem (3.1) with $\operatorname*{IC}_{s}^{\operatorname*{Green}}$ is well posed with a continuous solution operator denoted by $\operatorname*{S}_{\operatorname*{D}}\mathopen{}\mathclose{{}\left(s}\right):H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)\rightarrow H_{\operatorname*{loc}}^{1}(B_{R}^{+})$.
@@
      Then problem (3.1) with $\operatorname*{IC}_{0}^{\operatorname*{strong}}$ is well posed with a continuous solution operator denoted by ${\operatorname{S}_{\operatorname{D}}\mathopen{}\mathclose{{}\left(0}\right)}:H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)\rightarrow H_{\operatorname*{loc}}^{1}(B_{R}^{+})$.
@@
      Let $s=0$ and $n=2$. If and only if $R\neq 1$, the problem (3.1) with $\operatorname*{IC}_{0}^{\operatorname*{Green}}$ is well posed with a continuous solution operator denoted by ${\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}:H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)\rightarrow H_{\operatorname*{loc}}^{1}(B_{R}^{+})$.
@@
The results from Proposition 3.2 generalise to Dirichlet Helmholtz problems on the exterior of general bounded Lipschitz domains with the results in [12]. In this case, the condition $R\neq 1$ for the Laplace case in $n=2$ dimensions in Proposition 3.2.2b has to be replaced by the condition that the logarithmic capacity $\operatorname{Cap}_{\Gamma}$ defined in [12, p. 264] does not equal one, i.e., $\operatorname{Cap}_{\Gamma}\neq 1$.
@@ 3.2 The DtN operator for the Helmholtz problem on a sphere
The operator $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right):H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)\rightarrow H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right)$ depends continuously on the wavenumber $s\in\mathbb{C}^{\bullet}_{\geq 0}$ and admits a continuous extension to $\mathbb{C}_{\geq 0}$ with
@@
    $\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{R}}=\mathopen{}\mathclose{{}\left\langle g,\overline{\operatorname*{DtN}\mathopen{}\mathclose{{}\left(\overline{s}}\right)h}}\right\rangle_{S_{R}},$
@@
The proof of Theorem 3.4 in Subsection 4.4 employs an explicit representation of the $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)$ operator in terms of its Fourier coefficients derived in Section 4. The second main result is a Friedrichs-type inequality for functions with homogeneous DtN boundary conditions on $S_{R}$.
@@
Proof.
Recall $H^{1}(\Omega,\Gamma)$ for $\Gamma=S_{R}$ from (2.5) and consider the Hilbert space
$H=\mathopen{}\mathclose{{}\left\{v\in H^{1}\mathopen{}\mathclose{{}\left(\Omega,S_{R}}\right)\mid\partial_{r}v-\operatorname*{DtN}\mathopen{}\mathclose{{}\left(z}\right)v=0\ \text{ on }S_{R}\text{ for some }z\in\mathbb{C}_{\geq 0}}\right\}.$
(The completeness of $H$ follows from the continuity of the $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)$ operator by Theorem 3.4.) Assume for a contradiction to (3.6) that there exists a sequence $(v_{k})_{k\in\mathbb{N}}\subset H$ with $\mathopen{}\mathclose{{}\left\|v_{k}}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=1$ for all $k\in\mathbb{N}$ and $\|\nabla v_{k}\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}+\mathopen{}\mathclose{{}\left\|\partial_{r}v_{k}}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}\rightarrow 0$ as $k\rightarrow\infty$. Since this sequence is bounded in $H$, it admits a weakly convergent subsequence with weak limit $u\in H$. A standard compactness argument with the compact embedding $H^{1}\mathopen{}\mathclose{{}\left(\Omega}\right)\hookrightarrow L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)$ and the weak lower semicontinuity of $\|\nabla\bullet\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}$ reveal $\mathopen{}\mathclose{{}\left\|u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=1$ and $\mathopen{}\mathclose{{}\left\|\nabla u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=0$. Hence $u\in H$ is constant on the (in particular connected) domain $\Omega$. Because $u\in H$, there exists $s\in\mathbb{C}_{\geq 0}$ with $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)u=\partial_{r}u$. Since the constant $u$ has zero normal derivative, $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)u=0$ and Theorem 3.4.ii imply $\frac{n-2}{2}\mathopen{}\mathclose{{}\left\|u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(S_{R}}\right)}=0$. For $n\geq 3$, this shows $u=0$ and contradicts $\mathopen{}\mathclose{{}\left\|u}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}=1$. ∎
@@
$\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}\leq 2\sqrt{C_{\omega,1}^{-2}+1}\max\mathopen{}\mathclose{{}\left\{1,C_{\omega,1}C_{\operatorname*{F}}}\right\}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$
@@
Observe from (3.10) that
$\mathopen{}\mathclose{{}\left\|\Delta v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}\geq\mathopen{}\mathclose{{}\left(C_{\omega,1}C_{\operatorname*{F}}}\right)^{-2}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}-\mathopen{}\mathclose{{}\left(C_{\omega,1}^{-2}+1}\right)\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}^{2}$
and in turn
$\mathopen{}\mathclose{{}\left\|\Delta v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}\geq\mathopen{}\mathclose{{}\left(C_{\omega,1}C_{\operatorname*{F}}}\right)^{-1}\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}-\sqrt{C_{\omega,1}^{-2}+1}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$ (3.12)
Let $\max\mathopen{}\mathclose{{}\left\{1,C_{\omega,1}C_{\operatorname*{F}}}\right\}\mathopen{}\mathclose{{}\left|s}\right|^{2}\leq 1/2$ (which implies $\mathopen{}\mathclose{{}\left|s}\right|\leq 1$). Since $\Delta v=s^{2}v$, in $\omega$ it follows
@@
This allows for the estimate of the $L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)$ norm by
@@
(3.12)${\leq}\mathopen{}\mathclose{{}\left(1-\max\mathopen{}\mathclose{{}\left\{1,\mathopen{}\mathclose{{}\left(C_{\omega,1}C_{\operatorname*{F}}}\right)^{-1}}\right\}}\right)\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}+\max\mathopen{}\mathclose{{}\left\{1,C_{\omega,1}C_{\operatorname*{F}}}\right\}\sqrt{C_{\omega,1}^{-2}+1}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}$
$\leq\max\mathopen{}\mathclose{{}\left\{1,C_{\omega,1}C_{\operatorname*{F}}}\right\}\sqrt{C_{\omega,1}^{-2}+1}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$
It remains to prove the trace estimate in (3.8). An immediate consequence (known from [13, Cor. 3.2]) of a Young inequality and the trace estimate (2.4) reveals

### v2

@@ Abstract
This paper considers the Helmholtz problem in the exterior of a ball with Dirichlet boundary conditions and radiation conditions imposed at infinity. The differential Helmholtz operator depends on the complex wavenumber with non-negative real part and is formulated for general spatial dimensions. We prove wavenumber explicit continuity estimates of the corresponding Dirichlet-to-Neumann ($\operatorname*{DtN}$) operator which do not deteriorate as the complex wavenumber tends to zero.
@@ 1 Introduction
The theoretical justification and practical implementation of this approach require a thorough understanding of the involved DtN operator. In particular, for the numerical discretization of scattering problems, it is crucial to understand the explicit behaviour with respect to the wavenumber of the arising operators. Although the Fredholm theory in combination with (variants of) the unique continuation principle is a potent tool to establish well-posedness of acoustic problems, wavenumber explicit estimates require more refined analytical techniques. Among them are subtle definiteness (sign) properties of the real and imaginary part of the DtN operator [7, 9, 14, 18]. However, these properties have only been established for purely imaginary wavenumbers in $n=2,3$ dimensions [2, 14, 15], which prevents a unified theory for general complex wavenumbers: Existing results on Helmholtz problem based on the DtN approach are typically restricted to purely imaginary wavenumbers for $n\in\{2,3\}$, see, e.g., [7, 18] for recent developments on general coefficient Helmholtz problems.
To close this gap, this paper presents a unified mathematical analysis of spherical DtN operators for Helmholtz problems with general wavenumbers in any spatial dimension and provides a clearer mathematical understanding of the important definiteness characteristics of the DtN operator. In particular, our results enable a unified analysis of generalised boundary layer operators for Helmholtz problems with general coefficients and wavenumbers for $n\geq 2$ dimensions in [9]. The new insights also facilitate an improved tuning of control parameters in numerical methods depending on the real and imaginary part of the wavenumber such as [1] in future research.
@@
a) The wavenumber-dependent DtN operator admits an explicit Fourier series representation in terms of eigenfunctions of the Laplace-Beltrami operator and Bessel functions. A careful wavenumber-explicit control of the Fourier series coefficients enables sharp bounds on the real and imaginary part of the DtN operator for general wavenumbers and spatial dimensions. Our results extend the existing theory for purely imaginary wavenumbers, refining the bounds in [14, Sec. 3] for two dimensions and recovering the classical results [15, Eqn. (2.6.23)–(2.6.24)] for three dimensions.
b) The solution space of the Helmholtz equation in the truncated domain with transparent ($\operatorname*{DtN}$) boundary conditions on the artificial boundary satisfies a novel Friedrichs-type inequality in $n\geq 3$ dimensions. This enables wavenumber-explicit Sobolev and trace estimates in the natural weighted norms which do not deteriorate for small complex wavenumbers with non-negative real part. These estimates are the key to an explicit characterisation of the natural weighted trace norm for Helmholtz problems on exterior domains in [8, Thm. 3.4].
@@
Section 3 presents the exterior Helmholtz problem with Dirichlet boundary conditions and discusses the popular choices of the radiation conditions towards infinity. The well-posedness of the resulting exterior problems is established by Proposition 3.2. Its proof follows mainly the theory developed in [13] and is postponed to the appendix. Theorem 3.4 states the important definiteness properties of the DtN operator with sharp bounds on its real and imaginary parts. The Friedrichs-type inequality described in b) is formulated in Theorem 3.5 and leads in Corollary 3.6 to the announced wavenumber-explicit Sobolev and trace estimates.
Section 4 introduces series expansions of the $\operatorname*{DtN}$ operators for the different radiation conditions. Sharp wavenumber-explicit estimates of the series coefficients for non-zero wavenumbers are formulated as Theorem 4.1 and proved in Subsection 4.2. The special case of wavenumber zero (i.e., the Laplacian) requires extra care and is analysed in Subsection 4.3. This paper concludes with the proofs of Theorems 3.4 and 3.5 in Subsection 4.4.
@@ 2 Preliminaries
The physical problem is formulated in $\mathbb{R}^{n}$ for general dimension $n\in\mathopen{}\mathclose{{}\left\{2,3,\ldots}\right\}$. We consider Lipschitz domains $\omega\subset\mathbb{R}^{n}$ which are open and connected (not necessarily bounded) sets whose boundary is compact and locally the graph of a Lipschitz function [13, Def. 3.28]. The classical Lebesgue and Sobolev spaces on $\omega$ are denoted by $L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)$ and $H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ with norms $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}$ and $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)}$. For $\kappa\geq 0$, the space $H_{\operatorname*{loc}}^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ is given by all distributions $v$ on the compactly supported (relative to $\mathbb{R}^{n}$) and smooth functions $C_{\operatorname*{comp}}^{\infty}\mathopen{}\mathclose{{}\left(\omega}\right)$ such that $\varphi v\in H^{\kappa}\mathopen{}\mathclose{{}\left(\omega}\right)$ for all $\varphi\in C_{\operatorname*{comp}}^{\infty}\mathopen{}\mathclose{{}\left(\omega}\right)$. For the definition of Sobolev spaces $H^{\kappa}\mathopen{}\mathclose{{}\left(\partial\omega}\right)$ on $\mathopen{}\mathclose{{}\left(n-1}\right)$-dimensional boundaries $\partial\omega$ and their norm $\mathopen{}\mathclose{{}\left\|\cdot}\right\|_{H^{\kappa}\mathopen{}\mathclose{{}\left(\Gamma}\right)}$ we refer, e.g., to [13, pp. 96–99]. We define wavenumber-indexed norms [^1] For $s=0$, the quantity $\mathopen{}\mathclose{{}\left\|v}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right),0}=\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\omega}\right)}$ is only a seminorm in $H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)$. for $s\in\mathbb{C}_{\geq 0}$ by
@@
It is well-known that the outer unit normal $n(x)\in\mathbb{S}_{1}$ of the Lipschitz domain $\omega\subset\mathbb{R}^{n}$ exists for almost every $x\in\partial\omega$. The classical normal derivative $\partial/\partial n:\overline{\omega}\rightarrow\Gamma$ on some relatively open connectivity component $\Gamma\subset\partial\omega$ of the boundary can be continuously extended to a continuous mapping $\gamma_{1}:H^{1}\mathopen{}\mathclose{{}\left(\omega,\Delta}\right)\rightarrow H^{-1/2}\mathopen{}\mathclose{{}\left(\Gamma}\right)$ with
@@
which is equipped with the natural norm
$\mathopen{}\mathclose{{}\left\|\bullet}\right\|_{H^{1}(\omega,\Gamma)}:=\sqrt{\mathopen{}\mathclose{{}\left\|\bullet}\right\|_{H^{1}\mathopen{}\mathclose{{}\left(\omega}\right)}^{2}+\mathopen{}\mathclose{{}\left\|\gamma_{1}\bullet}\right\|_{H^{-1/2}(\Gamma)}^{2}}.$
The space $L(H_{1};H_{2})$ of continuous linear maps $A:H_{1}\to H_{2}$ between Hilbert spaces $H_{1}$ and $H_{2}$ is equipped with the operator norm topology. For the Fréchet space $H^{1}_{\mathrm{loc}}(\omega)=H_{2}$ as the image space, we set
$L(H_{1};H^{1}_{\mathrm{loc}}(\omega)):=\{A:H_{1}\to H^{1}_{\mathrm{loc}}(\omega)\ |\ \varphi A\in L(H_{1};H^{1}(\omega))\text{ for all }\varphi\in C^{\infty}_{\mathrm{comp}}(\omega)\}.$
Let $B_{R}\subset\mathbb{R}^{n}$ denote the open ball of radius $R>0$ at the origin. The outer unit normal vector $x/\mathopen{}\mathclose{{}\left\|x}\right\|$ on the boundary $S_{R}=\partial B_{R}$ points into the unbounded complement $B_{R}^{+}:=\mathbb{R}^{n}\setminus\overline{B_{R}}$ of $B_{R}$ and its derivative in this direction is denoted by $\partial_{r}$. The notation ||$\bullet$|| ∙ | is context-sensitive and may refer to the Lebesgue measure $|\omega|$ of a bounded measurable $n$-dimensional set $\omega\subset\mathbb{R}^{n}$, the surface measure $|\Gamma|$ of an $\mathopen{}\mathclose{{}\left(n-1}\right)$-dimensional manifold $\Gamma\subset\mathbb{R}^{n}$, and the cardinality $|J|$ of a countable set $J$.
@@ 3.1 The Dirichlet Helmholtz problem on the exterior of a ball
$\mathopen{}\mathclose{{}\left\{\begin{array}[c]{rl}\mathopen{}\mathclose{{}\left|u\mathopen{}\mathclose{{}\left(r\xi}\right)}\right|&\leq Cr^{\mu}\\ \mathopen{}\mathclose{{}\left|\frac{\partial u}{\partial r}\mathopen{}\mathclose{{}\left(r\xi}\right)+su\mathopen{}\mathclose{{}\left(r\xi}\right)}\right|&\leq Cr^{\mu-1}\end{array}}\right\}\quad\text{as }r\rightarrow\infty\text{ uniformly in }\xi\in S_{1}.$ (3.2a)
@@
In the following proposition we state the well-posedness of these exterior Dirichlet problems. We say that problem (3.1) is well posed for $\operatorname*{IC}_{s}$, if the corresponding solution operator $\operatorname{S_{D}}\in L(H^{1/2}(S_{R});H^{1}_{\mathrm{loc}}(B_{R}^{+}))$ that assigns each Dirichlet data $g\in H^{1/2}(S_{R})$ the unique solution $\operatorname{S_{D}}g\in H^{1}_{\mathrm{loc}}(B_{R}^{+})$ with $\operatorname{IC}_{s}$ to (3.1), is well-defined and continuous.
@@
      Problem (3.1) with $\operatorname*{IC}_{s}^{\operatorname*{Green}}$ is well posed with a continuous solution operator denoted by $\operatorname*{S}_{\operatorname*{D}}\mathopen{}\mathclose{{}\left(s}\right)\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H_{\operatorname*{loc}}^{1}(B_{R}^{+}))$.
@@
      Then problem (3.1) with $\operatorname*{IC}_{0}^{\operatorname*{strong}}$ is well posed with a continuous solution operator denoted by ${\operatorname{S}_{\operatorname{D}}\mathopen{}\mathclose{{}\left(0}\right)}\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H_{\operatorname*{loc}}^{1}(B_{R}^{+}))$.
@@
      Let $s=0$ and $n=2$. If and only if $R\neq 1$, the problem (3.1) with $\operatorname*{IC}_{0}^{\operatorname*{Green}}$ is well posed with a continuous solution operator denoted by ${\operatorname{S}^{\operatorname{log}}_{\operatorname{D}}}\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H_{\operatorname*{loc}}^{1}(B_{R}^{+}))$.
@@
The results from Proposition 3.2 generalise to Dirichlet Helmholtz problems on the exterior of general bounded Lipschitz domains with the results in [13]. In this case, the condition $R\neq 1$ for the Laplace case in $n=2$ dimensions in Proposition 3.2.2b has to be replaced by the condition that the logarithmic capacity $\operatorname{Cap}_{\Gamma}$ defined in [13, p. 264] does not equal one, i.e., $\operatorname{Cap}_{\Gamma}\neq 1$.
@@ 3.2 The DtN operator for the Helmholtz problem on a sphere
The first main result states the key properties of the DtN operator including bounds on its real and imaginary parts. The symmetry and the boundedness, given here in terms of the natural wavenumber-indexed norms, are more standard and included with elementary proofs for completeness.
@@
The operator $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\in L(H^{1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right);H^{-1/2}\mathopen{}\mathclose{{}\left(S_{R}}\right))$ depends continuously on the wavenumber $s\in\mathbb{C}^{\bullet}_{\geq 0}$ and admits a continuous extension to $\mathbb{C}_{\geq 0}$ with
@@
    $\mathopen{}\mathclose{{}\left\langle\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)g,\overline{h}}\right\rangle_{S_{R}}=\mathopen{}\mathclose{{}\left\langle g,\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)\overline{h}}\right\rangle_{S_{R}}=\mathopen{}\mathclose{{}\left\langle g,\overline{\operatorname*{DtN}\mathopen{}\mathclose{{}\left(\overline{s}}\right)h}}\right\rangle_{S_{R}},$
@@
The second main result is a Friedrichs-type inequality for functions with homogeneous DtN boundary conditions on $S_{R}$ in $n\geq 3$ dimensions. The proofs of Theorems 3.4 and 3.5 follow in Subsection 4.4 below based an explicit representation of the $\operatorname*{DtN}\mathopen{}\mathclose{{}\left(s}\right)$ operator derived in Section 4 in terms of its Fourier coefficients.
@@
$\mathopen{}\mathclose{{}\left\|v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}\leq 2C_{\operatorname{F}}\sqrt{1+C_{\omega,1}^{2}}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$ (3.9)
@@
Let $\max\mathopen{}\mathclose{{}\left\{1,C_{\omega,1}C_{\operatorname*{F}}}\right\}\mathopen{}\mathclose{{}\left|s}\right|^{2}\leq 1/2$ (which implies $\mathopen{}\mathclose{{}\left|s}\right|\leq 1$) and observe from $\Delta v=s^{2}v$ in $\omega$ that
@@
This and (3.10) establish the estimate (3.9) of the $L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)$ norm by
@@
$\leq C_{\operatorname*{F}}\sqrt{1+C_{\omega,1}^{2}}\mathopen{}\mathclose{{}\left\|\nabla v}\right\|_{L^{2}\mathopen{}\mathclose{{}\left(\Omega}\right)}.$
It remains to prove the trace estimate in (3.8). An immediate consequence (known from [14, Cor. 3.2]) of a Young inequality and the trace estimate (2.4) reveals